5.) Find the three measure of central tendency(the mean, median, and mode).
7,11,17,18,22

Answer by nerdybill(7384)

Find the three measure of central tendency(the mean, median, and mode).
7,11,17,18,22
Mean:
(7+11+17+18+22)/5 = 75/5 = 15
Median: (middle value)
17
Mode: (number occurring most frequent)
No mode